# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

### v2

@@ 2.1. Algorithmic fairness: A gentle introduction
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). In order to apply the CFR, rights need to be balanced (Article 52(1) CFR). First, legislators must balance different CFR rights proportionally when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this balancing obligation. A logic of proportionality of the different rights guides Fundamental Rights assessments (Almada and Petit, 2023). The Fundamental Rights are not absolute, hierarchical, or quantifiable and need to be applied case by case (Sousa e Silva, 2024). This balancing and proportionality can make it very difficult to grasp the exact concept of Fundamental Rights.
Further complications arise due to the fact that the provisions of the CFR target public parties in relation to private parties (vertical applicability) (Fornasier, 2015). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (Fornasier, 2015; Cherednychenko, 2007; Frantziou, 2015; Prechal, 2020) and in the AI Act specifically (Lewis et al., 2025).
2.3. Interaction of law and computer science: Differentiating between concepts.
Law and computer science are different fields. While this becomes clear from the different viewpoints above, it also leads to severe challenges when discussing algorithmic fairness and law: the terminology differs. The difficulty stems from the fact that different fields and society at certain times and places define fairness and discrimination in different ways (Sousa e Silva, 2024). While the compatibility of fairness metrics with normative legal statements has already been discussed (Deck et al., 2024a; Hellman, 2020; Wachter et al., 2021, 2020), there is still an inherent difficulty in combining these two fields. This interdisciplinary challenge needs to acknowledge that computer science fairness and legal non-discrimination regulation are moving in the same direction; however, they present different concepts (Deck et al., 2024b). Although inspired by law, the technical algorithmic fairness debate primarily focuses on detecting, preventing and mitigation unfairness in algorithms(Deck et al., 2024b), while legal non-discrimination law focuses on normative statements.
As shown in the previous section, the algorithmic fairness domain uses terms like bias or fairness. These concepts are difficult to grasp and are topics of debate regarding their exact meaning (Sousa e Silva, 2024; Genovesi et al., 2024; Xenidis and Senden, 2020; Bosoer et al., 2023). Furthermore, in legal terms, algorithmic fairness is rarely used.
When we discuss algorithmic fairness and non-discrimination regulation, we are going to differentiate terms. First, (algorithmic) fairness refers to the AI-related domain of analyzing, evaluating, auditing, and mitigating unequal treatment between (groups of) individuals based on normative notions (Laux et al., 2024; Estévez Almenzar et al., 2022). When we use the terms (non-)discrimination, we will refer to classical non-discrimination law in the European Union.
3. EU AI Act’s non-discrimination regulations.
3.1. The EU AI Act: History, Scope, and Definitions
Emergent technology needs efficient regulation. Before the work on the European AI Act started, the EU High-Level Expert Group on Artificial Intelligence had already developed ethical guidelines for AI systems. These guidelines were later further developed for the AI Act (Koumeri et al., 2023). While the AI Act is not the first AI regulation worldwide (Wodi, 2024), it presents harmonized rules and obligations for the use and “placing on the market” (Article 1(2) AIA) of AI systems in the European Union. Thus, a level playing field for AI products within the Union is established. The AI Act is (predominantly) a product safety law (Almada and Petit, 2023). In fact, the original drafts did not have any individual rights components (Hacker et al., 2024b). The Members of the European Parliament did not prioritize non-discrimination regulation in the beginning (Chiappetta, 2023).
Risk is a central concept within the AI Act (Hacker et al., 2024a). Similar to other digital regulations in the EU, the AI Act defines risk in Article 3(2) AIA. A risk “means the combination of the probability of an occurrence of harm and the severity of that harm”. Based on this definition, the AI Act differentiates between different risk types with special requirements, which is not uncontroversial (Bosoer et al., 2023): prohibited systems (Article 5 AIA), high-risk systems (Article 6 ff. AIA), certain AI systems (Article 50 AIA), and GPAI models (Article 51 ff. AIA). Prohibited systems pose an unacceptable risk (Nolte et al., 2024). In contrast, high-risk systems pose a risk that can be handled through extensive regulations. Certain AI Systems are systems that can imitate human content, are emotion recognition systems, or biometric categorization systems, and therefore users must be informed transparently about their use. GPAI Models are, in computer science terms, (Multimodal) Foundation models, which are a separate category due to their capability to serve various purposes, both for direct use and integration in other AI systems (Article 3(66) AIA). If a system does not fall into one of these categories, only a limited set of rules applies. One example is the obligation that providers and deployers of AI systems ensure the AI literacy of their staff or other persons dealing with the system (Article 4 AIA).
@@ 3.1. The EU AI Act: History, Scope, and Definitions
The AI Act has specific wording for the difference between model and system. On the one hand, the AI Act defines in Article 3(1) AIA an AI system as: “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments”. While the AIA sticks to the AI system definition when introducing high-risk systems, it refers to GPAI models in Article 51 ff. AIA. For GPAI Model Rec. 97 AIA defines that “Although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as for example a user interface, to become AI systems.” From a computer science perspective, it is unusual not to use the term “model” or “machine learning model”, which are the terms most often used in technical papers. Since we are discussing legal norms, we stick to definitions provided by the AI Act, see also (Nolte et al., 2024).
@@
Providers (Article 3(3) AIA) and deployers (Article 3(4) AIA) are the central addressees of the AI Act. Providers are the entities that develop and place a model on the market or put the system into service. In contrast, the deployers are using an AI system under its authority. Many of the following obligations focus on the providers of a system (Article 16 AIA) compared to the deployers (Article 26 AIA).
3.2. Non-discrimination regulations within the AI Act: From clarity to vagueness.
@@
Analyzing non-discrimination regulations starts with identifying regulations. In total, we scanned the AI Act for non-discrimination-related terms (from specific to unspecific concepts): discrimination, Fundamental Rights, fairness, and bias. We noticed that within the definitions of Article 3 AIA, the serious incident (Article 3(49) AIA) and systemic risk (Article 3(65) AIA) refer to Fundamental Rights. Thus, these were included in our analysis as well.
@@
While analyzing the regulations, we observed that terms appear in the majority of cases in the recitals and only very limited in the articles of the AI Act. The recitals are only considered helpful when interpreting the articles (Klimas and Vaiciukaite, 2008) but do not have legal binding force (noa, 2014). Therefore, our analysis does not focus on the recitals.
Table 1 presents our main results. Not all articles directly relate to specific AI regulations but sometimes regulate market authorities, for example, Article 28 AIA. Thus, we focused on core AI regulation and present the list of articles mentioning non-discrimination-related terms in Table 1. A full table with all articles of the AI Act including non-discrimination-related terms, can be found in the appendix in Table 2.
Article
Discrimination /
discriminatory
Fundamental
Rights
Fair /
fairness
Bias
Systemic
risk
Serious
incident
Specific Topic
General Introduction
1 Subject matter of AI Act
2 Scope of the AI Act
3 Definitions
High-risk Systems and its obligations
6 Classification rules for high-risk AI systems
7 Amendments to Annex III AI Act
9 Risk management system for High-Risk AI Systems
10 Data and data governance for High-Risk AI Systems
13 Transparency and provision of information
14 Human oversight of High-risk Systems
15 Accuracy, robustness and cybersecurity
17 Quality management system
26 Obligations of deployers of high-risk AI systems
27 Fundamental Rights impact assessment for high-risk AI systems
GPAI Models
51 Classification of AI models with systemic risk
52 Notification Procedure for GPAI models
53 Obligations for providers of general-purpose AI models
54 Authorised representatives of providers of GPAI models
55 Obligations of providers of GPAI Models with systemic risk
56 Codes of practice for GPAI Models
Table 1. Non-discrimination and algorithmic fairness-related terms in the AI Act for selected most important articles. A full overview is provided in 2 in the appendix. From left to right, regulatory clarity decreases.
@@ 56 Codes of practice for GPAI Models
This analysis shows, first, that the majority of non-discrimination regulations in the EU AI Act concern the regulation of high-risk systems. GPAI models are only regulated implicitly by the systemic risk and serious incident terms. This will have an impact on our further analysis: we focus on high-risk systems compared to GPAI models.
@@
Second, most non-discrimination regulations are only implicitly addressed by the reference to Fundamental Rights. Discrimination and bias as threats are explicitly addressed only for high-risk systems. But here too, we find a relation to Fundamental Rights.
@@
Third, the term “fairness” from algorithmic fairness is not used at all within the AI Act. On the one hand, this result is surprising. The fairness aspect reveals a gap between wording in the computer science and ML community and the legal field. On the other hand, however, as the non-discrimination rules are based on established rules within EU legislative frameworks, it is not a surprising finding.
@@
After analyzing which parts of the AI Act contain non-discrimination regulations, we will now analyze the details for high-risk systems and general purpose AI systems in detail.
3.3. High-risk AI non-discrimination regulations.
@@
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino, 2024) than other less risky systems. It is expected that 5%-15% [^10] For a critical assessment of these numbers see (Almada and Petit, 2023) of all AI systems fall under the category of high-risk systems (Commission, 2021).
@@
We identified Articles 9, 10, and 15 AIA as main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^11] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law. All other Articles listed in tab 1 will be discussed and summarized in the last paragraph.
@@
Article 9 AIA: Undefined discrimination regulation within risk management systems Article 9 of the AI Act covers risk management systems of High-Risk Systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI System. All risks after mitigation, called residual risks, need to be judged acceptable (Soler Garrido et al., 2023).
@@
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or Fundamental Rights. As explained in Section 2.2, Fundamental Rights also include the right of non-discrimination. Thus, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues. Zuiderveen Borgesius et al. (2024) pointed out that this Article mandates the identification and mitigation of discrimination in AI systems. The initial draft of the AI Act did not include the reference to Fundamental Rights, and it was added by the EU Council in 2022 121212see Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022. (Schuett, 2024), highlighting the need to ensure non-discriminatory regulations.
@@
We identified three main problematic issues with the non-discrimination regulation in Article 9 AIA. First, only risks that are known and reasonably foreseeable need to be identified and analyzed. However, this is particularly problematic for Large Multimodal Models (LMMs) in high-risk contexts (Hacker et al., 2024a). What constitutes a reasonably foreseeable risk in an era where large language models are still being developed and not fully understood? Although alignment techniques have developed (Ji et al., 2023), discrimination still persists in LMMs (Gupta et al., 2023). When are these risks foreseeable? The AI Act leaves too much room for unclear speculation. From a technical perspective, benchmarks to judge the risk associated with LLMs have been proposed (Guldimann et al., 2024); however, automated bias testing is far from trivial, and we currently lack adequate ML techniques. Recital 65 AIA clarifies that the most appropriate risk-management measures should be used. There is a danger that even the most appropriate measures are not sufficient to identify foreseeable risks.
@@
Second, the link to Fundamental Rights leaves many potential non-discrimination factors unaddressed (Kusche, 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited by Article 21(1) CFR but may be prohibited by other European or national laws. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the Fundamental Rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche, 2024).
@@
This issue is not unique to Article 9(2) AIA but is present in all Articles of the AI Act referring to Fundamental Rights. Health and safety are “bread and butter of product safety law” (Almada and Petit, 2023). The added link to Fundamental Rights adds a different dimension compared to product safety requirements. This mixture of aims is problematic since product safety law and Fundamental Rights law have different target regimes: technical aspects vs. individual rights.
@@
Third, the relation between known and reasonably foreseeable risks and direct and indirect discrimination in traditional EU non-discrimination law is not well explored. Even if we assume that the risks through Fundamental Rights are indirectly related to non-discrimination requirements, it is unclear what known and reasonably foreseeable implies. Referring back to Section 2.2, traditional discrimination law differentiates between direct and indirect discrimination. Is (only) direct discrimination foreseeable? While indirect discrimination through algorithms seems more plausible, is it foreseeable? The exact relation is undefined.
@@
Article 10(2)(f) AIA: Main input non-discrimination regulation of high-risk system. The core of the non-discrimination regulation for high-risk systems is Article 10 AIA. Article 10(2)(f) AIA requires that training, validation, and testing datasets shall be subject to an “examination in view of possible biases that are likely to affect the health and safety of persons, have a negative impact on fundamental rights or lead to discrimination prohibited under Union law, especially where data outputs influence inputs for future operations”.
@@
Bias is a vague term. The first observation we make is that the input data of an AI system should be subjected to a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of what bias means (see also Section 2.1) (van Bekkum, 2024). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (Hacker et al., 2024b). This makes it very hard to determine the regulatory content.
@@
Do the negative effects need to be certain or likely? It is an open question to which condition the wording “that are likely” refers to. If the wording only refers to the first condition, then the negative impact on Fundamental Rights has to be certain. If the wording “that are likely” refers to the Fundamental Rights or to discrimination prohibited under Union law, then these conditions also include a probability statement. When comparing the wording of Article 10(2)(f) AIA to Article 9(2)(a) AIA, we note that Article 9(2)(a) targets the risk of Fundamental Rights impact. The risk is balancing between the likelihood of an occurrence and the severity of harm (Article 3(2) AIA). It therefore seems more consistent to assume that the term “likely” in Article 10(2)(f) AIA refers to all conditions. This view is also supported by the translation of the AI Act in different languages such as German, where it is more evident that the wording “that are likely” applies to all conditions.
@@
Article 10(2)(f) AIA contrasts with Article 9(2)(a) AIA. In contrast to Article 10 AIA, the bias must have a likely negative impact on Fundamental Rights or lead to discrimination, while Article 9 uses the word “risk”. Again, risk is a combination of likelihood and severity. Compared to the risk in Article 9(2) AIA, in Article 10(2)(f) AIA, the severity of the harm is not taken into account, only the likelihood. This difference between Article 9(2) AIA and 10(2)(f) AIA creates unnecessary regulatory uncertainty. It remains unclear why the legislator opted for the difference in applicability between Article 9 AIA and Article 10 AIA.
@@
Through the wording of Article 10(2)(f) AIA, the regulation directly links to traditional non-discrimination law. This condition covers many different aspects of discriminatory effects. In general, this achieves a strong protection objective. The relationship between direct and indirect discrimination remains unclear. Is indirect discrimination likely? Furthermore, intersectional aspects of discrimination are only vaguely covered by traditional discrimination law (see Sec. 2.2). Are they likely as well? When artificial intelligence systems use many different features as input, intersectional discrimination is likely (Wang et al., 2022; Roy et al., 2023).

### v3

@@ 2 A primer on non-discrimination from an AI and a legal perspective
However, the right to non-discrimination is not absolute. It is important to note that the right to non-discrimination is only one right in the CFR, among others, such as the freedom of expression and information (Article 11(1) CFR). In order to apply the CFR, rights need to be balanced (Article 52(1) CFR). First, legislators must balance different CFR rights proportionally when creating new laws. Additionally, when laws reference the CFR, implementers must also follow this balancing obligation. A logic of proportionality of the different rights guides fundamental rights assessments (Almada and Petit 2023). Fundamental rights are not absolute, hierarchical, or quantifiable and need to be applied case by case (Sousa e Silva 2024). The provisions of the CFR target public parties (vertical applicability)(Fornasier 2015). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (Fornasier 2015; Cherednychenko 2007; Frantziou 2015; Prechal 2020) and in the AI Act context specifically (Lewis et al. 2025). In the context of the GDPR, scholars and the European Court of Justice (ECJ) noted that the CFR right can have a horizontal effect if the secondary law reflects a general principle of EU law (Ufert 2020). Since the AI Act also reflects general principles of EU law, we do not exclude some horizontal applicability of the CFR. When discussing fundamental rights in the AI Act, it is also important to keep the political and institutional landscape in mind (Palmiotto 2025).
Direct and indirect discrimination in EU Law. While the CFR is part of the primary law, it also influences the law of the European Union (Lewis et al. 2025). EU law has a long history of non-discrimination regulations and different laws include non-discrimination regulations [^7] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination˙en.. While also primary law include non-discrimnation regulations, one example is the prohibition of unequal payment based on gender (Article 157(1) TFEU), the focus of traditional non-discrimnation law in the EU is on secondary law (directives or regulations). For example, the Directive 2000/43/EC (Race Equality Directive) and Directive 2000/78/EC (Employment Equality Directive) include specific non-discrimination regulations.
In order to legally assess discrimination under the specific secondary EU Law, traditionally two different versions of discrimination are differentiated: direct and indirect discrimination (Zuiderveen Borgesius et al. 2024; Wachter, Mittelstadt, and Russell 2020). However, importantly, EU-based direct discrimination does not require any intentional wrongdoing (Weerts et al. 2023; Xenidis and Senden 2020; Adams-Prassl, Binns, and Kelly-Lyth 2023).
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against by using a (listed) protected attribute.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (Zuiderveen Borgesius et al. 2024).
Also, a third category of discrimination needs to be taken into account [^8] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may only occur at the intersection of gender and age(Weerts et al. 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al. 2023; Xenidis 2023; Atrey 2018).
3 EU AI Act’s non-discrimination regulations for high-risk systems.
Before we discuss the specific non-discrimination regulations for high-risk systems, we will briefly introduce the history, scope, and most important definitions of the AI Act. This short introduction is followed by an overview of the AI Act’s non-discrimination regulations for high-risk systems.
The EU AI Act: History, Scope, and Definitions
In this section, we first describe the origins of the EU AI Act and clarify key concepts such as risk, systems, and the difference between developers and deployers, which are necessary to understand our analysis of the AI Act.
Emergent technology can benefit from efficient regulation. As one of the first comprehensive AI regulations worldwide (Wodi 2024), the AI Act (AIA) introduces harmonized rules and obligations for the use and “placing on the market” of AI systems within the European Union (Article 1(2) AIA). It is primarily a product safety regulation (Almada and Petit 2023), aiming to establish a level playing field for AI technologies across the Union (Nolte, Rateike, and Finck 2024). Notably, the original drafts of the EU Commission did not include explicit provisions on individual rights (Hacker et al. 2024). At this stage, individual rights — understood as protecting the rights of individuals affected — were largely absent. In fact, Members of the European Parliament initially did not prioritize the regulation of algorithmic discrimination (Chiappetta 2023). However, over the course of the legislative process, different aspects of non-discrimination regulations were integrated into the final text.
@@
The AI Act regulates AI systems and also includes provisions for General Purpose AI (GPAI) models. The relationship or distinction between a model and a system in the context of the AIA remains legally unresolved. Article 3(1) defines an “AI system” as “a machine-based system that is designed to operate with varying levels of autonomy and that may exhibit adaptiveness after deployment, and that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments“, but the article does not provide a definition of “AI model”. Recital 97 clarifies that “although AI models are essential components of AI systems, they do not constitute AI systems on their own. AI models require the addition of further components, such as, for example, a user interface, to become AI systems.” This formulation does not conclusively settle the terminological relationship between models and systems, as discussed in more detail in Nolte, Rateike, and Finck (2024). Moreover, the regulation of “AI systems” contrasts with the (trustworthy) computer science literature, which typically studies machine learning models. Since the focus here is on the legal framework, we adopt the terminology used in the AI Act.
The core regulatory structure of the AI Act is built on a risk-based approach (Hacker, Cordes, and Rochon 2024). Risk is defined in Article 3(2) AIA as “the combination of the probability of the occurrence of harm and the severity of that harm.” Based on this definition, the AI Act categorizes AI systems into different risk levels, each associated with a specific set of regulatory requirements [^9] This classification has been subject to criticism, see e.g., Bosoer, Cantero Gamito, and Rubio-Marin (2023).:
  • Unacceptable risk (Art. 5 AIA): These AI systems are prohibited. Examples include social scoring by governments.
  • High-risk (Art. 6ff. AIA): These systems are subject to stringent obligations, such as requirements for robustness, accuracy, or non-discrimination. Examples include AI used in recruitment, credit scoring, or law enforcement.
  • Certain AI systems (Art. 50 AIA) with specific risk: These systems must meet transparency obligations, such as informing users they are interacting with an AI system. Examples include systems that produce deep-fakes.
  • All other systems: These systems are not subject to specific regulatory obligations under the AIA, except broad regulations like Art. 4 AIA (AI Literacy).
In addition, the AIA includes specific provisions for General Purpose AI (GPAI) models (Articles 51 ff. AIA). The use case is not predefined, which is specific to GPAI models. LLMs are an example of a classical GPAI model. The AI Act distinguishes between general-purpose models posing “systemic risk” and those that do not.
@@
The primary addressees of the AI Act are providers (Article 3(3) AIA) and deployers (Article 3(4) AIA). Providers are the entities responsible for developing and placing an AI model on the market or putting an AI system into service. In contrast, deployers are those who use an AI system under their authority. Many of the obligations we discuss in the following sections primarily concern the providers of AI systems, while some apply specifically to deployers (e.g., Article 26 AIA).
Non-discrimination regulations within the AI Act.
@@
Analyzing non-discrimination regulations starts with identifying regulations. In total, we scanned the AI Act for non-discrimination-related terms: discrimination, fundamental right, fairness, and bias. We noticed that within the definitions of Article 3 AIA, the terms serious incident (Article 3(49) AIA) and systemic risk (Article 3(65) AIA) refer to fundamental rights. Thus, these were included in our analysis as well.
@@
While analyzing the regulations, we observed that terms appear in the majority of cases in the recitals, rather than in the articles of the AI Act. The recitals are only considered helpful when interpreting the articles (Klimas and Vaiciukaite 2008) but do not have a legally binding force. Therefore, our analysis does not focus on the recitals. A full table with all articles of the AI Act, including non-discrimination-related terms, can be found in the appendix in Table 1.
@@
This analysis shows, first, that the majority of non-discrimination regulations in the EU AI Act concern the regulation of high-risk systems. GPAI models are only implicitly regulated by the systemic risk and serious incident terms. This will have an impact on our further analysis: we focus on high-risk systems compared to GPAI models.
@@
Second, most non-discrimination regulations are only implicitly addressed by the reference to fundamental rights. Discrimination and bias as threats are explicitly addressed only for high-risk systems. Here too, we observe a relation to fundamental rights.
@@
Third, the term “fairness” from algorithmic fairness is not used at all within the AI Act. On the one hand, this result is surprising. The absence of the fairness aspect reveals a gap between the wording used in the computer science and ML communities and that of the legal field. On the other hand, however, as the non-discrimination rules are based on established rules within EU legislative frameworks, it is not a surprising finding.
@@
After analyzing which parts of the AI Act contain non-discrimination regulations, we will now analyze the details for high-risk systems in detail.
High-risk AI non-discrimination regulations.
@@
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino 2024) than other less risky systems. Recital 48 states that the adverse impact caused by the AI system on fundamental rights is of particular relevance when classifying an AI system as a high-risk system. It is expected that 5%-15% [^10] For a critical assessment of these numbers see (Almada and Petit 2023) of all AI systems fall under the category of high-risk systems (Commission 2021).
@@
We identified Articles 9, 10, and 15 AIA as the main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^11] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law.. Furthermore, Article 11 (together with Annex III) and Article 13 AIA will be discussed due to their relationship to the aforementioned articles.
@@
Article 9 AIA: Non-discrimination regulations within risk management systems Article 9 of the AI Act covers risk management systems of high-risk systems. According to Article 9(2)(a-d) AIA, risk management systems are designed to identify, evaluate, and mitigate risks of AI systems in a continuous, iterative process throughout the entire lifecycle of a high-risk AI system. All risks after mitigation, called residual risks, need to be judged acceptable (Soler Garrido et al. 2023).
@@
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or fundamental rights. Through the link to fundamental rights, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues (Zuiderveen Borgesius et al. 2024). The initial draft of the AI Act did not include the reference to fundamental rights, and it was added by the EU Council in 2022  [^12] See Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022..
@@
We identified three challenges from an algorithmic fairness perspective with the non-discrimination regulation in Article 9 AIA. First, only risks that are known and reasonably foreseeable need to be identified and analyzed. However, this is particularly problematic for Large Multimodal Models (LMMs) in high-risk contexts (Hacker, Cordes, and Rochon 2024). What constitutes a reasonably foreseeable risk in an era where large language models are still being developed and not fully understood? Although alignment techniques have been developed (Ji et al. 2023), discrimination still persists in LMMs (Gupta et al. 2023). When are these risks foreseeable? The AI Act leaves much room for speculation. From a technical perspective, benchmarks to judge the risk associated with LLMs have been proposed (Guldimann et al. 2024); however, automated bias testing is far from trivial, and we currently lack adequate ML techniques. Recital 65 AIA clarifies that the most appropriate risk-management measures should be used. There is a danger that even the most appropriate measures are not sufficient to identify foreseeable risks.
@@
Second, the link to fundamental rights leaves potential non-discrimination factors unaddressed (Kusche 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited by Article 21(1) CFR but may be prohibited by other European or national laws. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the fundamental rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche 2024).
@@
This issue is not unique to Article 9(2) AIA but is present in all articles of the AI Act referring to fundamental rights. Health and safety are “bread and butter of product safety law” (Almada and Petit 2023). The added link to fundamental rights adds a different dimension compared to product safety requirements. This mixture of aims is problematic since product safety law and fundamental rights law have different target regimes: technical aspects vs. individual rights.
@@
Third, the relation between known and reasonably foreseeable risks and direct and indirect discrimination in traditional EU non-discrimination law is not well explored (Zuiderveen Borgesius et al. 2024; Wachter, Mittelstadt, and Russell 2020). Even if we assume that the risks through fundamental rights are indirectly related to non-discrimination requirements, it is unclear what known and reasonably foreseeable implies. Is (only) direct discrimination foreseeable? While indirect discrimination through algorithms seems more plausible, is it foreseeable? The exact relation is undefined.
@@
Article 10(2)(f) AIA: Main input non-discrimination regulation of high-risk system. The core of the non-discrimination regulation for high-risk systems is Article 10(2)(f) AIA. Article 10(2)(f) AIA requires that training, validation, and testing datasets shall be subject to an “examination in view of possible biases that are likely to affect the health and safety of persons, have a negative impact on fundamental rights or lead to discrimination prohibited under Union law, especially where data outputs influence inputs for future operations”.
@@
The first observation we make is that the input data of an AI system should be subjected to a bias analysis. Most importantly, the term bias is neither defined in the AI Act nor is there a common understanding of what bias means (van Bekkum 2024). It seems that the regulators had a more technical definition of bias in mind, focusing on the diversity of training data in different dimensions compared to social, ethical, or structural biases (Hacker et al. 2024). This could imply difficulties in determining the regulatory content.
@@
Do the negative effects need to be certain or likely? It is an open question to which condition the wording “that are likely” refers to. If the wording only refers to the first condition, then the negative impact on fundamental rights has to be certain. If the wording “that are likely” refers to the fundamental rights or to discrimination prohibited under Union law, then these conditions also include a probability statement. When comparing the wording of Article 10(2)(f) AIA to Article 9(2)(a) AIA, we note that Article 9(2)(a) targets the risk of fundamental rights impact. The “risk” is a balance between the likelihood of an occurrence and the severity of harm (Article 3(2) AIA). It therefore seems more consistent to assume that the term “likely” in Article 10(2)(f) AIA refers to all conditions. This view is also supported by the translation of the AI Act in different languages such as German, where it is more evident that the wording “that are likely” applies to all conditions.
@@
Article 10(2)(f) AIA contrasts with Article 9(2)(a) AIA. In contrast to Article 10 AIA, the bias must have a likely negative impact on fundamental rights or lead to discrimination, while Article 9 uses the word “risk”. Again, risk is a combination of likelihood and severity. Compared to the risk in Article 9(2) AIA, in Article 10(2)(f) AIA, the severity of the harm is not taken into account, only the likelihood. This difference between Article 9(2) AIA and 10(2)(f) AIA creates regulatory uncertainty. It remains open why the legislator opted for the difference in applicability between Article 9 AIA and Article 10 AIA.
@@
Through the wording of Article 10(2)(f) AIA, the regulation directly links to traditional non-discrimination law. Many different aspects of discriminatory effects are covered. On the one hand, the link achieves a strong protection objective for the input side of the AI systems. However, even without this link, Union law, which prohibits discrimination, could be applicable (see also next section). Thus, the question remains: what new legislative content is added? On the other hand, all existing problems within traditional non-discrimination law remain. For example, the relationship between direct and indirect discrimination remains unclear. Is indirect discrimination likely? Furthermore, intersectional aspects of discrimination are only vaguely covered by traditional discrimination law. Are they likely as well? When artificial intelligence systems use many different features as input, intersectional discrimination is likely (Wang, Ramaswamy, and Russakovsky 2022; Roy, Horstmann, and Ntoutsi 2023).